THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 116 — 2 April 2012

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Figure 1: Supernova 1987A resolved for the first time at 3 mm wavelength, offering a preview of what ALMA will show. See Lakićević, Zanardo, et al. (2012).

Editorial

Dear Colleagues,

It is my pleasure to present you the 116^{th} issue of the Magellanic Clouds Newsletter. There are fascinating new results on the supernova remnants SN 1987A, B 0509−67.5 and B 0519−69.0, and a lot of new work on Cepheid variables, plus a variety of other very interesting new findings about the Magellanic clouds and their internal processes and content.

The next issue is planned to be distributed on the 1^{st} of June 2012.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

An absence of ex-companion stars in the Type Ia supernova remnant SNR 0509−67.5

Bradley E. Schaefer^{1} and Ashley Pagnotta^{1}

1Louisiana State University, USA

Type Ia supernova (SNe Ia) are thought to originate in the explosion of a white dwarf. The explosion could be triggered by the merger of two white dwarfs (’double-degenerate’ origin), or by mass transfer from a companion star (the ’single-degenerate’ path). The identity of the progenitor is still controversial; for example, a recent argument against the single-degenerate origin has been widely rejected. One way to distinguish between the double- and singledegenerate progenitors is to look at the center of a known SN Ia remnant to see whether any former companion star is present. A likely ex-companion star for the progenitor of Tycho’s supernova has been identified, but that claim is still controversial. Here we report that the central region of the supernova remnant SNR 0509−67.5 (the site of a Type Ia supernova 400 ± 50 years ago, based on its light echo) in the Large Magellanic Cloud contains no ex-companion star to a limit of V = 26.9 magnitude (M_{V} = +8.4) within the extreme 99.73% region with radius 1.′′43. The lack of any ex-companion star to deep limits rules out all published single-degenerate models. The only remaining possibility is that the progenitor for this particular SN Ia was a double-degenerate system.

Published in Nature, 481, 164 (2012) Available from arXiv:1201.2195

The progenitor of the Type Ia supernova that created SNR 0519−69.0 in the Large Magellanic Cloud

Zachary I. Edwards^{1}, Ashley Pagnotta^{2} and Bradley E. Schaefer^{2}

1Columbus State University, Columbus, GA 31907, USA 2Louisiana State University, Baton Rouge, LA 70803, USA

Models for the progenitor systems of Type Ia supernovae can be divided into double-degenerate systems, which contain two white dwarfs, and single-degenerate systems, which contain one white dwarf plus one companion star (either a red giant, a subgiant, or a > 1.16 M_{⊙} main sequence star). The white dwarf is destroyed in the supernova explosion, but any non-degenerate companion remains intact. We present the results of a search for an ex-companion star in SNR 0519−69.0, located in the Large Magellanic Cloud, based on images taken with the Hubble Space Telescope with a limiting magnitude of V = 26.05. SNR 0519−69.0 is confidently known to be from a Type Ia supernova based on its light echoes and X-ray spectra. The geometric center of the remnant (based on the Hα and X-ray shell) is at 05:19:34.83, −69:02:06.92 (J2000). Accounting for the measurement uncertainties, the orbital velocity, and the kick velocity, any ex-companion star must be within 4.′′7 of this position at the 99.73% confidence level. This circle contains 27 main sequence stars brighter than V = 22.7, any one of which could be the ex-companion star left over from a supersoft source progenitor system. The circle contains no post-main sequence stars, and this rules out the possibility of all other published single-degenerate progenitor classes (including symbiotic stars, recurrent novae, helium donors, and the spin-up/spin-down models) for this particular supernova. The only remaining possibility is that SNR 0519−69.0 was formed from either a supersoft source or a double-degenerate progenitor system.

Accepted for publication in ApJ Letters Available from arXiv:1201.6377

On the application of Wesenheit function in deriving distance to Galactic Cepheids

Chow-Choong Ngeow^{1}

1NCU, Taiwan

In this work, we explore the possibility of using theWesenheit function to derive individual distances to Galactic Cepheids, as the dispersion of the reddening free Wesenheit function is smaller than the optical period luminosity (P–L) relation. When compared to the distances from various methods, the averaged differences between our results and published distances range from −0.061 to 0.009, suggesting that the Wesenheit function can be used to derive individual Cepheid distances. We have also constructed Galactic P–L relations and selected Wesenheit functions based on the derived distances. A by-product from this work is the derivation of Large Magellanic Cloud distance modulus when calibrating the Wesenheit function. It is found to be 18.531 ± 0.043 mag.

Published in The Astrophysical Journal Available from arXiv:1202.0339

Period–Luminosity relations for Cepheid variables: From mid-infrared to multi-phase

Chow-Choong Ngeow^{1}, Shashi Kanbur^{2}, Earl P. Bellinger^{2}, Marcella Marconi^{3}, Ilaria Musella^{3}, Michele Cignoni^{4} and Ya-Hong Lin^{1}

1NCU, Taiwan 2SUNY-Oswego, USA 3Osservatorio Astronomico di Capodimonte, Italy 4Bologna University, Italy

This paper discusses two aspects of current research on the Cepheid period-luminosity (P–L) relation: the derivation of mid-infrared (MIR) P–L relations and the investigation of multi-phase P–L relations. The MIR P–L relations for Cepheids are important in the James Webb Space Telescope era for the distance scale issue, as the relations have potential to derive the Hubble constant within ∼ 2% accuracy – a critical constraint in precision cosmology. Consequently, we have derived the MIR P–L relations for Cepheids in the Large and Small Magellanic Clouds, using archival data from Spitzer Space Telescope. We also compared currently empirical P–L relations for Cepheids in the Magellanic Clouds to the synthetic MIR P–L relations derived from pulsational models. For the study of multi-phase P–L relations, we present convincing evidence that the Cepheid P–L relations in the Magellanic Clouds are highly dynamic quantities that vary significantly when considered as a function of pulsational phase. We found that there is a difference in P–L relations as a function of phase between the Cepheids in each of the Clouds; the most likely cause for this is the metallicity difference between the two galaxies. We also investigated the dispersion of the multi-phase P–L relations, and found that the minimum dispersion do not differ significantly from the mean light P–L dispersion.

Accepted for publication in Astrophysics and Space Science Available from arXiv:1202.0150

Anomalous Cepheids in the Large Magellanic Cloud: Insight on their origin and connection with the star formation history

G. Fiorentino^{1} and M. Monelli^{2}

1INAF–Osservatorio Astronomico di Bologna, via Ranzani 1, 40127, Bologna, Italy 2Instituto de Astrofísica de Canarias, Calle Via Lactea s/n, E38205 La Laguna, Tenerife, Spain

The properties of variable stars can give independent constraints on the star formation history of the host galaxy, through the determination of the age and the metallicity of the parent population. We investigate the pulsation properties of 84 Anomalous Cepheids (ACs) detected by the OGLE-III survey in the Large Magellanic Cloud (LMC), in order to understand the formation mechanism and the characteristics of the parent population they originated from. We used updated theoretical pulsation scenario to derive the mass and the pulsation mode of each AC in the sample. We also discuss, by means of a Kolmogorov–Smirnov test, the spatial distribution of the ACs, in comparison with that of other groups of variable stars, and connect their properties with the star formation history of the LMC. We find that the ACs’ mean mass is 1.2 ± 0.2 M_{⊙}. We show that ACs do not follow the same spatial distribution of classical Cepheids. This, and the fact that their period–luminosity relations are different, provides further support to the hypothesis that ACs are not the extension to low luminosity of classical Cepheids. The spatial distribution of ACs is also different from that of bona-fide tracers of the old population, such as RR Lyrae stars and population II Cepheids. We therefore suggest that the majority of ACs in the LMC is made of intermediate-age (1–6,Gyr), metalpoor single stars. Finally, we investigate the relation between the frequency of ACs and the luminosity of the host galaxy, disclosing that purely old system follow a very tight relation, and that galaxies with important intermediateage and young star formation tend to have an excess of ACs, in agreement with the fact that they host ACs formed via both single and binary stars channels.

Accepted for publication in A&A Available from arXiv:1202.2752

On the evolutionary and pulsation mass of Classical Cepheids: III. The case of the eclipsing binary Cepheid CEP 0227 in the Large Magellanic

Cloud

P.G. Prada Moroni^{1,2}, M. Gennaro^{3}, G. Bono^{4,5}, G. Pietrzyński^{6,7}, W. Gieren^{6}, B. Pilecki^{6,7}, D. Graczyk^{6} and I.B. Thompson^{8}

1Dipartimento di Fisica, Università di Pisa, Largo B. Pontecorvo 2, 56127 Pisa, Italy 2INFN–Pisa, via E. Fermi 2, 56127 Pisa, Italy 3Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117, Heidelberg, Germany 4Dipartimento di Fisica, Università di Roma Tor Vergata, Via della Ricerca Scientifica 1, 00133 Roma, Italy 5INAF–Osservatorio Astronomico di Roma, Via Frascati 33, 00040 Monte Porzio Catone, Italy 6Universidad de Concepción, Departamento de Astronomia, Casilla 160-C, Concepción, Chile 7Obserwatorium Astronomiczne Uniwersytetu Warszawskiego, Aleje Ujazdowskie 4, 00-478 Warszawa, Poland 8Carnegie Observatories, 813 Santa Barbara Street, Pasadena, CA 91101-1292, USA

We present a new Bayesian approach to constrain the intrinsic parameters (stellar mass, age) of the eclipsing binary system CEP 0227 in the Large Magellanic Cloud. We computed several sets of evolutionary models covering a broad range in chemical compositions and in stellar mass. Independent sets of models were also constructed either by neglecting or by including a moderate convective core overshooting (β_{ov} = 0.2) during central hydrogen burning phases. Sets of models were also constructed either by neglecting or by assuming a canonical (η = 0.4, 0.8) or an enhanced (η = 4) mass-loss rate. The most probable solutions were computed in three different planes: luminosity– temperature, mass–radius and gravity–temperature. By using the Bayes Factor, we found that the most probable solutions were obtained in the gravity–temperature plane with a Gaussian mass prior distribution. The evolutionary models constructed by assuming a moderate convective core overshooting (β_{ov} = 0.2) and a canonical mass loss rate (η = 0.4) give stellar masses for the primary (Cepheid), M = 4.14_{−0.05}^{+0.04} M_{⊙}, and for the secondary, M = 4.15_{−0.05}^{+0.04} M_{⊙}, that agree at the 1% level with dynamical measurements. Moreover, we found ages for the two components and for the combined system, t = 151_{−3}^{+4} Myr, that agree at the 5% level. The solutions based on evolutionary models that neglect the mass loss attain similar parameters, while those ones based on models that either account for an enhanced mass loss or neglect convective core overshooting have lower Bayes Factors and larger confidence intervals. The dependence on the mass loss rate might be the consequence of the crude approximation we use to mimic this phenomenon. By using the isochrone of the most probable solution and a Gaussian prior on the LMC distance, we found a true distance modulus, 18.53_{−0.02}^{+0.02} mag, and a reddening value, E(B − V) = 0.142_{−0.010}^{+0.005} mag, that agree quite well with similar estimates in the literature.

Accepted for publication in ApJ Available from arXiv:1202.2855

The influences of the Magellanic Clouds on the Galaxy: Pole shift, warp, and star formation history

Kenji Bekki^{1}

1ICRAR (UWA), Australia

We investigate how the Large Magellanic Cloud (LMC) influences the evolution of the Galaxy after the LMC enters into the virial radius of the dark matter halo of the Galaxy for the first time. Both the Galaxy and the LMC are modeled as N-body particles in our models so that the dynamical influences of the LMC on the Galaxy can be investigated in a fully self-consistent manner. Furthermore, the orbital parameters for the LMC are carefully chosen such that the present location of the LMC in the Galaxy can be rather precisely reproduced in our simulations. We particularly investigate the influences of the LMC on the precession rate, the outer stellar and gaseous structures, and the star formation history of the Galaxy. Our principals results are summarized as follows. The LMC–Galaxy dynamical interaction can cause “pole shift” (or irregular precession/nutation) of the Galaxy and the typical rate of pole shift (θ̇_{d}) is ∼ 2^{∘} Gyr^{−1} corresponding to ∼ 7 μas yr^{−1}. The LMC–Galaxy interaction induces the formation of the outer warp structures of the Galaxy, which thus confirms the results of previous numerical simulations on the formation of the Galactic warp. The LMC–Galaxy interaction also induces the formation of outer (R > 20 kpc) spiral arms and increases the vertical velocity dispersion of the outer disk significantly. The mean star formation rate of the Galaxy for the last several Gyrs can be hardly influenced by the LMC’s tidal force. The age and metallicity distribution of stars in the solar-neighborhood (7 kpc <≤ R < 10 kpc) for the last several Gyr can be only slightly changed by the past LMC–Galaxy interaction. If the LMC is accreted onto the Galaxy as a group with small dwarf galaxies, then the stripped dwarfs forms a unique polar distribution within the Galaxy. Based on these results, we discuss how the possible ongoing Galactic pole shift with θ̇_{d} ∼ 10 μas yr^{−1} can be detected by future observational studies by Gaia.

Accepted for publication in MNRAS Available from arXiv:1202.0358

Comparison of convective overshooting models and their impact on abundances from integrated light spectroscopy of young (< 3 Gyr) star

clusters

J.E. Colucci^{1} and R.A. Bernstein^{1}

1University of California Santa Cruz, USA

As part of an ongoing program to measure detailed chemical abundances in nearby galaxies, we use a sample of young to intermediate age clusters in the Large Magellanic Cloud with ages of 10 Myr to 2 Gyr to evaluate the effect of isochrone parameters, specifically core convective overshooting, on Fe abundance results from high resolution, integrated light spectroscopy. In this work we also obtain fiducial Fe abundances from high resolution spectroscopy of the cluster individual member stars. We compare the Fe abundance results for the individual stars to the results from isochrones and integrated light spectroscopy to determine whether isochrones with convective overshooting should be used in our integrated light analysis of young to intermediate age (10 Myr –3 Gyr) star clusters. We find that when using the isochrones from the Teramo group, we obtain more accurate results for young and intermediate age clusters over the entire age range when using isochrones without convective overshooting. While convective overshooting is not the only uncertain aspect of stellar evolution, it is one of the most readily parametrized ingredients in stellar evolution models, and thus important to evaluate for the specific models used in our integrated light analysis. This work demonstrates that our method for integrated light spectroscopy of star clusters can provide unique tests for future constraints on stellar evolution models of young and intermediate age clusters.

Accepted for publication in Astrophysical Journal Available from arXiv:1202.1524

Yellow and red supergiants in the Large Magellanic Cloud

Kathryn F. Neugent^{1}, Philip Massey^{1}, Brian Skiff^{1} and Georges Meynet^{2}

1Lowell Observatory, USA 2Geneva University, Switzerland

Due to their transitionary nature, yellow supergiants provide a critical challenge for evolutionary modeling. Previous studies within M 31 and the SMC show that the Geneva evolutionary models do a poor job at predicting the lifetimes of these short-lived stars. Here we extend this study to the LMC while also investigating the galaxy’s red supergiant content. This task is complicated by contamination by Galactic foreground stars that color and magnitude criteria alone cannot weed out. Therefore, we use proper motions and the LMC’s large systemic radial velocity (∼ 278 km s^{−1}) to separate out these foreground dwarfs. After observing nearly 2,000 stars, we identified 317 probable yellow supergiants, 6 possible yellow supergiants and 505 probable red supergiants. Foreground contamination of our yellow supergiant sample was ∼ 80%, while that of the the red supergiant sample was only 3%. By placing the yellow supergiants on the H–R diagram and comparing them against the evolutionary tracks, we find that new Geneva evolutionary models do an exemplary job at predicting both the locations and the lifetimes of these transitory objects.

Accepted for publication in ApJ Available from arXiv:1202.4225

The tidal origin of the Magellanic Stream and the possibility of a stellar counterpart

Jonathan Diaz^{1,2} and Kenji Bekki^{1}

1ICRAR, University of Western Australia, Australia 2Institute of Astronomy, University of Cambridge, UK

We present an N-body model that reproduces the morphology and kinematics of the Magellanic Stream (MS), a vast neutral hydrogen (H i) structure that trails behind the Large and Small Magellanic Clouds (LMC and SMC, respectively) in their orbit about the Milky Way. After investigating 8 million possible orbits consistent with the latest proper motions, we adopt an orbital history in which the LMC and SMC have only recently become a strongly interacting binary pair. We find that their first close encounter 2 Gyr ago provides the necessary tidal forces to disrupt the disk of the SMC and thereby create the MS. The model also reproduces the on-sky bifurcation of the two filaments of the MS, and we suggest that a bound association with the Milky Way is required to reproduce the bifurcation. Additional H i structures are created during the tidal evolution of the SMC disk, including the Magellanic Bridge, the ”Counter Bridge”, and two branches of leading material. Insights into the chemical evolution of the LMC are also provided, as a substantial fraction of the material stripped away from the SMC is engulfed by the LMC. Lastly we compare three different N-body realizations of the stellar component of the SMC, which we model as a pressuresupported spheroid motivated by recent kinematical observations. We find that an extended spheroid is better able to explain the stellar periphery of the SMC, and the tidal evolution of the spheroid may imply the existence of a stellar stream akin to the gaseous MS.

Accepted for publication in ApJ Available from arXiv:1112.6191

The optical depth of H ii regions in the Magellanic Clouds

E.W. Pellegrini^{1}, M.S. Oey^{1}, P.F. Winkler^{2}, S.D. Points^{3} and R.C. Smith^{3}

1Department of Astronomy, University of Michigan, 500 Church Street, Ann Arbor, MI, 48109, USA 2Department of Physics, Middlebury College, Middlebury, VT 05753, USA 3Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatory, Casilla 603, La Serena, Chile

We exploit ionization-parameter mapping as a powerful tool to measure the optical depth of star-forming H ii regions. Our simulations based on the Cloudy photoionization code and our new, SurfBright surface brightness simulator demonstrate that this technique can directly diagnose most density-bounded, optically thin nebulae with spatially resolved emission line data. We apply this method to the Large and Small Magellanic Clouds, using the data from the Magellanic Clouds Emission Line Survey. We generate new H ii region catalogs based on photoionization criteria set by the observed ionization structure in the [S ii]/[O iii] ratio and Hα surface brightness. The luminosity functions from these catalogs generally agree with those from Hα-only surveys. We then use ionization-parameter mapping to crudely classify all the nebulae into optically thick vs. optically thin categories, yielding fundamental new insights into the Lyman continuum radiation transfer. We find that in both galaxies, the frequency of optically thin objects correlates with Hα luminosity, and that the numbers of these objects dominate above Ł(Hα) ≥ 10^{37.0}. Similarly, the frequency of optically thick regions correlates with H i column density, with optically thin objects dominating at the lowest N(H i). The integrated escape luminosity of ionizing radiation is dominated by the largest regions, and corresponds to luminosity-weighted, ionizing escape fractions from the H ii region population of ≥ 0.42 and ≥ 0.44 in the LMC and SMC, respectively. These values correspond to global galactic escape fractions of 4% and 11%, respectively. This is sufficient to power the ionization rate of the observed diffuse ionized gas in both galaxies. Since our optical depth estimates tend to be underestimates, and also omit the contribution from field stars without nebulae, our results suggest the possibility of significant galactic escape fractions of Lyman continuum radiation.

Submitted to Astrophysical Journal Available from arXiv:1202.3334

Preliminary results for RR Lyrae stars and Classical Cepheids from the VISTA Magellanic Cloud (VMC) survey

V. Ripepi^{1}, M.I. Moretti^{2,3}, G. Clementini^{3}, M.R. Cioni^{4,5}, J.B. Marquette^{6} and P. Tisserand^{7}

1INAF–Osservatorio Astronomico di Capodimonte, Via Moiariello 16, 80131, Naples, Italy 2INAF–Osservatorio Astronomico di Bologna, via Ranzani 7, Bologna, Italy 3University of Bologna, Department of Astronomy, via Ranzani 7, 40127, Bologna, Italy 4University of Hertfordshire, Physics Astronomy and Mathematics, Hatfield AL10 9AB, UK 5University Observatory Munich, Scheinerstraße 1, 81679 München, Germany – Research Fellow of the Alexander von Humboldt Foundation 6UPMC–CNRS, UMR7095, Institut d’Astrophysique de Paris, F-75014, Paris, France 7Research School of Astronomy & Astrophysics, Mount Stromlo Observatory, Cotter Road, Weston ACT 2611, Australia

The VISTA Magellanic Cloud (VMC, P.I.: M.R. Cioni) survey is collecting K_{s}-band time series photometry of the system formed by the two Magellanic Clouds (MC) and the “bridge” that connects them. These data are used to build K_{s}-band light curves of the MC RR Lyrae stars and Classical Cepheids and determine absolute distances and the 3D geometry of the whole system using the K-band period–luminosity (P LK_{s}), the period–luminosity–color (P LC) and the Wesenheit relations applicable to these types of variables. As an example of the survey potential we present results from the VMC observations of two fields centered respectively on the South Ecliptic Pole and the 30 Doradus star forming region of the Large Magellanic Cloud. The VMC K_{s}-band light curves of the RR Lyrae stars in these two regions have very good photometric quality with typical errors for the individual data points in the range of ∼ 0.02 to 0.05 mag. The Cepheids have excellent light curves (typical errors of ∼ 0.01 mag). The average K_{s} magnitudes derived for both types of variables were used to derive P LK_{s} relations that are in general good agreement within the errors with the literature data, and show a smaller scatter than previous studies.

Accepted for publication in Ap&SS Available from arXiv:1202.5863

Testing the effects of opacity and the chemical mixture on the excitation of pulsations in B stars of the Magellanic Clouds

S. Salmon^{1}, J. Montalbán^{1}, T. Morel^{1}, A. Miglio^{1,2}, M.-A. Dupret^{1} and A. Noels^{1}

1Institut d’Astrophysique et de Géophysique de l’Université de Liège, Allée du 6 Août 17, 4000 Liège, Belgium 2School of Physics and Astonomy, University of Birmingham, Edgbaston, Birmingham B15 2TT, UK

The B-type pulsators known as β Cephei and Slowly Pulsating B (SPB) stars present pulsations driven by the κ mechanism, which operates thanks to an opacity bump due to the iron group elements. In low-metallicity environments such as the Magellanic Clouds, β Cep and SPB pulsations are not expected. Nevertheless, recent observations show evidence for the presence of B-type pulsator candidates in both galaxies. We seek an explanation for the excitation of β Cep and SPB modes in those galaxies by examining basic input physics in stellar modelling: i) the specific metal mixture of B-type stars in the Magellanic Clouds; ii) the role of a potential underestimation of stellar opacities. We first derive the present-day chemical mixtures of B-type stars in the Magellanic Clouds. Then, we compute stellar models for that metal mixture and perform a non-adiabatic analysis of these models. In a second approach, we simulate parametric enhancements of stellar opacities due to different iron group elements. We then study their effects in models of B stars and their stability. We find that adopting a representative chemical mixture of B stars in the Small Magellanic Cloud cannot explain the presence of B-type pulsators there. An increase of the opacity in the region of the iron-group bump could drive B-type pulsations, but only if this increase occurs at the temperature corresponding to the maximum contribution of Ni to this opacity bump. We recommend an accurate computation of Ni opacity to understand B-type pulsators in the Small Magellanic Cloud, as well as the frequency domain observed in some Galactic hybrid β Cep-SPB stars.

Accepted for publication in MNRAS Available from arXiv:1203.0527

Characterizing cosmic-ray propagation in massive star-forming regions: The case of 30 Doradus and the Large Magellanic Cloud

Eric J. Murphy^{1}, Troy A. Porter^{2,3}, Igor V. Moskalenko^{2,3}, George Helou^{4} and Andrew W. Strong^{5}

1Observatories of the Carnegie Institution for Science, 813 Santa Barbara Street, Pasadena, CA 91101, USA 2Hansen Experimental Physics Laboratory, Stanford University, Stanford, CA 94305, USA 3Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305, USA 4California Institute of Technology, MC 314-6, Pasadena, CA 91125, USA 5Max-Planck-Institut für extraterrestrische Physik, Postfach 1312, D-85741 Garching bei München, Germany

Using infrared, radio, and γ-ray data, we investigate the propagation characteristics of cosmic-ray (CR) electrons and nuclei in the 30 Doradus (30 Dor) star-forming region in the Large Magellanic Cloud (LMC) using a phenomenological model based on the radio–far-infrared correlation within galaxies. Employing a correlation analysis, we derive an average propagation length of ∼ 100–140 pc for ∼ 3 GeV CR electrons resident in 30 Dor from consideration of the radio and infrared data. Assuming that the observed γ-ray emission towards 30 Dor is associated with the star-forming region, and applying the same methodology to the infrared and γ-ray data, we estimate a ∼ 20 GeV propagation length of 200–320 pc for the CR nuclei. This is approximately twice as large as for ∼ 3 GeV CR electrons, corresponding to a spatial diffusion coefficient that is ∼ 4 times higher, scaling as (R/GV)^{δ} with δ ≈ 0.7–0.8 depending on the smearing kernel used in the correlation analysis. This value is in agreement with the results found by extending the correlation analysis to include ∼ 70 GeV CR nuclei traced by the 3–10 GeV γ-ray data (δ ≈ 0.66 ± 0.23). Using the mean age of the stellar populations in 30 Dor and the results from our correlation analysis, we estimate a diffusion coefficient D_{R} ≈ 0.9–1.0 × 10^{27}(R/GV)^{0.7} cm^{2} s^{−1}. We compare the values of the CR electron propagation length and surface brightness for 30 Dor and the LMC as a whole with those of entire disk galaxies. We find that the trend of decreasing average CR propagation distance with increasing disk-averaged star formation activity holds for the LMC, and extends down to single star-forming regions, at least for the case of 30 Dor.

Accepted for publication in Astrophysical Journal Available from arXiv:1203.1626

The life and death of dense molecular clumps in the Large Magellanic Cloud

Jonathan P. Seale^{1,2}, Leslie W. Looney^{1}, Tony Wong^{1}, Juergen Ott^{3}, Uli Klein^{4} and Jorge L. Pineda^{5}

1University of Illinois, USA 2Space Telescope Science Institute, USA 3National Radio Astronomy Observatory, USA 4Universität Bonn Argelander-Institut für Astronomie, Germany 5Jet Propulsion Laboratory, USA

We report the results of a high spatial (parsec) resolution HCO^{+} (J = 1–0) and HCN (J = 1–0) emission survey toward the giant molecular clouds of the star formation regions N 105, N 113, N 159, and N 44 in the Large Magellanic Cloud. The HCO^{+} and HCN observations at 89.2 and 88.6 GHz, respectively, were conducted in the compact configuration of the Australia Telescope Compact Array. The emission is imaged into individual clumps with masses between 10^{2} and 10^{4} M_{⊙} and radii of < 1 pc to ∼ 2 pc. Many of the clumps are coincident with indicators of current massive star formation, indicating that many of the clumps are associated with deeply-embedded forming stars and star clusters. We find that massive YSO-bearing clumps tend to be larger (> 1 pc), more massive (M > 10^{3} M_{⊙}), and have higher surface densities (∼ 1 g cm^{−2}), while clumps without signs of star formation are smaller (< 1 pc), less massive (M < 10^{3} M_{⊙}), and have lower surface densities (∼ 0.1 g cm^{−2}). The dearth of massive (M > 10^{3} M_{⊙}) clumps not bearing massive YSOs suggests the onset of star formation occurs rapidly once the clump has attained physical properties favorable to massive star formation. Using a large sample of LMC massive YSO mid-IR spectra, we estimate that ∼ 2/3 of the massive YSOs for which there are Spitzer mid-IR spectra are no longer located in molecular clumps; we estimate that these young stars/clusters have destroyed their natal clumps on a time scale of at least 3 × 10^{5} yrs.

Accepted for publication in ApJ Available from arXiv:1203.4505

Post-AGB stars in the SMC as tracers of stellar evolution: the extreme s-process enrichment of the 21 μm star J 004441.04−732136.4

Kenneth De Smedt^{1}, Hans Van Winckel^{1}, Amanda I. Karakas^{2}, Lionel Siess^{3}, Stephane Goriely^{3} and Peter R. Wood^{2}

1Instituut voor Sterrenkunde, K.U. Leuven, Celestijnenlaan 200D, B-3001, Belgium 2Research School of Astronomy and Astrophysics, Mount Stromlo Observatory, Weston Creek ACT 2611, Australia 3Institut d’Astronomie et d’Astrophysique, Université Libre de Bruxelles, ULB, CP 226, 1050 Brussels, Belgium

Context: This paper is part of a larger project in which we want to focus on the still poorly understood AGB third dredge-up processes and associated s-process nucleosynthesis. Aims: We confront accurate spectral abundance analyses of post-AGB stars in both the LMC and SMC, to state-of-the-art AGB model predictions. With this comparison we aim at improving our understanding of the 3^{rd} dredge-up phenomena and their dependencies on initial mass and metallicity. Methods: Because of the well constrained distance with respect to Galactic post-AGB stars, we choose an extragalactic post-AGB star for this contribution, namely the only known 21 μm object of the Small Magellanic Cloud (SMC): J 004441.04−732136.4. We used optical UVES spectra to perform an accurate spectral abundance analysis. With photometric data of multiple catalogues we construct a spectral energy distribution and perform a variability analysis. The results are then compared to predictions of tailored theoretical chemical AGB evolutionary models for which we used two evolution codes. Results: Spectral abundance results reveal J 004441.04−732136.4 to be one of the most s-process enriched objects found up to date, while the photospheric C/O ratio of 1.9±0.7, shows the star is only modestly C-rich. J 004441.04−732136.4 also displays a low [Fe/H] = −1.34 ± 0.32, which is significantly lower than the mean metallicity of the SMC. From the SED, a luminosity of 7600 ± 200 L_{⊙} is found, together with E(B − V) = 0.64 ± 0.02. According to evolutionary post-AGB tracks, the initial mass should be approximately 1.3 M_{⊙} . The photometric variability shows a clear period of 97.6 ± 0.3 days. The detected C/O as well as the high s-process overabundances (e.g., [Y/Fe] = 2.15, [La/Fe] = 2.84) are hard to reconcile with the predictions. The chemical models also predict a high Pb abundance, which is not compatible with the detected spectrum, and a very high ^{12}C/^{13}C, which is not yet constrained by observations. The predictions are only marginally dependent on the evolution codes used. Conclusions: By virtue of their spectral types, favourable bolometric corrections as well as their constrained distances, post-AGB stars in external galaxies offer unprecedented tests to AGB nucleosynthesis and dredge-up predictions. We focus here on one object J 004441.04−732136.4, which is the only known 21 μm source of the SMC. We show that our theoretical predictions match the s-process distribution, but fail in reproducing the detected high overabundances and predict a high Pb abundance which is not detected. Additionally, there remain serious problems in explaining the observed pulsational properties of this source.

Accepted for publication in A&A Available from arXiv:1203.4413

Zooming in on Supernova 1987A at sub-mm wavelengths

Maša Lakićević^{1,2}, Jacco Th. van Loon ^{2}, Thomas Stanke^{1}, Carlos De Breuck^{1} and Ferdinando Patat^{1}

1ESO, Germany 2Keele University, UK

Supernova 1987A (SN 1987A) in the neighbouring Large Magellanic Cloud offers a superb opportunity to follow the evolution of a supernova and its remnant in unprecedented detail. Recently, far-infrared (far-IR) and sub-mm emission was detected from the direction of SN 1987A, which was interpreted as due to the emission from dust, possibly freshly synthesized in the SN ejecta. To better constrain the location and hence origin of the far-IR and submm emission in SN 1987A, we have attempted to resolve the object in that part of the electro-magnetic spectrum. We observed SN 1987A during July–September 2011 with the Atacama Pathfinder EXperiment (APEX), at a wavelength of 350 μm with the Submillimetre APEX Bolometer CAmera (SABOCA) and at 870 μm with the Large APEX BOlometer CAmera (LABOCA). The 350-μm image has superior angular resolution (8′′) over that of the Herschel Space Observatory 350-μm image (25′′). The 870-μm observation (at 20′′ resolution) is a repetition of a similar observation made in 2007. In both images, at 350 and 870 μm, emission is detected from SN 1987A, and the source is unresolved. The flux densities in the new (2011) measurements are consistent with those measured before with Herschel at 350 μm (in 2010) and with APEX at 870 μm (in 2007). A higher dust temperature (≈ 33 K) and lower dust mass might be possible than what was previously thought. The new measurements, at the highest angular resolution achieved so far at far-IR and sub-mm wavelengths, strengthen the constraints on the location of the emission, which is thought to be close to the site of SN 1987A and its circumstellar ring structures. These measurements set the stage for upcoming observations at even higher angular resolution with the Atacama Large Millimeter Array (ALMA).

Accepted for publication in A&A Letters Available from arXiv:1203.4975

The remnant of Supernova 1987A resolved at 3 mm wavelength

Maša Lakićević^{1,2}, Giovanna Zanardo^{3}, Jacco Th. van Loon^{2}, Lister Staveley-Smith^{3,5}, Toby Potter^{3}, C.-Y. Ng^{4} and B.M. Gaensler^{5,6}

1ESO, Germany 2Keele University, UK 3ICRAR, University of Western Australia, Australia 4Department of Physics, McGill University, Montréal, Canada 5ARC Centre of Excellence for All-sky Astrophysics (CAASTRO), Australia 6Sydney Institute for Astronomy (SIfA), Australia

The proximity of core-collapse Supernova 1987A (SN 1987A) in the Large Magellanic Cloud (LMC) and its rapid evolution make it a unique case study of the development of a young supernova remnant. We aim at resolving the remnant of SN 1987A for the first time in the 3-mm band (at 94 GHz). We observed the source at 3-mm wavelength with a 750-m configuration of the Australia Telescope Compact Array (ATCA). We compare the image with a recent 3-cm image and with archival X-ray images. We present a diffraction-limited image with a resolution of 0.′′7, revealing the ring structure seen at lower frequencies and at other wavebands. The emission peaks in the eastern part of the ring. The 3-mm image bears resemblance to early X-ray images (from 1999–2000). We place an upper limit of 1 mJy (2 σ) on any discrete source of emission in the centre (inside of the ring). The integrated flux density at 3 mm has doubled over the six years since the previous observations at 3 mm. At 3 mm – i.e. within the operational domain of the Atacama Large Millimeter/submillimeter Array (ALMA) – SN 1987A appears to be dominated by synchrotron radiation from the inner rim of the equatorial ring, characterised by moderately-weak shocks. There is no clear sign of emission of a different nature, but the current limits do not rule out such component altogether.

Accepted for publication in A&A Letters Available from arXiv:1203.4982

The Optical Gravitational Lensing Experiment. The catalog of stellar proper motions toward Magellanic Clouds

R. Poleski^{1}, I. Soszyński^{1}, A. Udalski^{1}, M. K. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, Ł. Wyrzykowski^{1,3} and K. Ulaczyk^{1}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2Universidad de Concepción, Departamento de Astronomia, Casilla 160-C, Concepción, Chile 3Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK

We present a catalog of over 6.2 million stars with measured proper motions. All these stars are observed in the direction of the Magellanic Clouds within the brightness range 12 < I < 19 mag. Based on these proper motions around 440 000 Galactic foreground stars can be selected. Because the proper motions are based on a few hundred epochs collected during eight years, their statistical uncertainties are below 0.5 mas yr^{−1} for stars brighter than I = 18.5 mag. The parallaxes are derived with uncertainties down to 1.6 mas. For above 13 000 objects parallaxes are derived with significance above 3σ, which allows selecting around 270 white dwarfs (WDs). The search for common proper motion binaries among stars presented was performed resulting in over 500 candidate systems. The most interesting ones are candidate halo main sequence star–WD and WD–WD systems. The application of the catalog to empirically bound the Cepheid instability strip is also discussed.

Submitted to Acta Astronomica Available from arXiv:1203.2649 and from ftp://ftp.astrouw.edu.pl/ogle/ogle3/pm/mcs/

Conference Paper

A kinematic study of the giant star-forming region 30 Doradus

Sergio Torres-Flores^{1}, Rodolfo Barbá^{1,2}, Jesús Maíz Apellániz^{3}, Mónica Rubio^{4} and Guillermo Bosch^{5}

1Departamento de Física, Universidad de La Serena, Chile 2ICATE–CONICET, San Juan, Argentina 3Instituto de Astrofísica de Andalucía–CSIC, Spain 4Departamento de Astronomía, Universidad de Chile, Chile 5Facultad de Ciencias Astronómicas y Geofísicas, UNLP, Argentina

We present, for the first time, an optical spectroscopic data cube of the giant star-forming region 30 Doradus, obtained with the GIRAFFE on the VLT at Paranal Observatory. The main emission lines present in this data cube correspond to Hα, [N ii] 6548 Å and [N ii] 6584 Å. By using this data set, we found that Hα presents from simple to multiple profiles, which suggests that different physical mechanisms act in different ways on the excited gas in 30 Doradus. We found, at least, three unclassified large expanding structures. These structures correlate with peaks in the Xray distribution. Given the excellent signal-to-noise ratio and the large spatial coverage of this data cube, we have studied in detail the kinematics of 30 Doradus, showing the importance of the small scale phenomena on the integrated properties of 30 Doradus.

Oral contribution, published in 54^{th} Bulletin of the Argentine Astronomical Society (BAAA 54) Available from arXiv:1202.2036

Review Papers

Asymptotic Giant Branch variables in the Galaxy and the Local Group

Patricia Whitelock^{1,2}

1SAAO, South Africa 2UCT, South Africa

AGB variables, particularly the large amplitude Mira type, are a vital step on the distance scale ladder. They will prove particularly important in the era of space telescopes and extremely large ground-based telescopes with adaptive optics, which will be optimized for infrared observing. Our current understanding of the distances to these stars is reviewed with particular emphasis on improvements that came from Hipparcos as well as on recent work on Local Group galaxies. In addition to providing the essential calibration for extragalactic distances Gaia may also provide unprecedented insight into the poorly understood mass-loss process itself.

Published in Astrophysics and Space Science Available from arXiv:1201.2997 and from http://www.springerlink.com/content/x5073u16422p7827/

Low-mass pre–main-sequence stars in the Magellanic Clouds

Dimitrios A. Gouliermis^{1}

1Max Planck Institute for Astronomy, Königstuhl 17, 69117 Heidelberg, Germany

The stellar Initial Mass Function (IMF) suggests that stars with sub-solar mass form in very large numbers. Most attractive places for catching low-mass star formation in the act are young stellar clusters and associations, still (half- )embedded in star-forming regions. The low-mass stars in such regions are still in their pre–main-sequence (PMS) evolutionary phase, i.e. they have not started their lives on the main-sequence yet. The peculiar nature of these objects and the contamination of their samples by the fore- and background evolved populations of the Galactic disk impose demanding observational techniques, such as X-ray surveying and optical spectroscopy of large samples for the detection of complete numbers of PMS stars in the Milky Way. The Magellanic Clouds, the metal-poor companion galaxies to our own, demonstrate an exceptional star formation activity. The low extinction and stellar field contamination in star-forming regions of these galaxies imply a more efficient detection of low-mass PMS stars than in the Milky Way, but their distance from us make the application of the above techniques unfeasible. Nonetheless, imaging with the Hubble Space Telescope within the last five years yield the discovery of solar and sub-solar PMS stars in the Magellanic Clouds from photometry alone. Unprecedented numbers of such objects are identified as the low-mass stellar content of star-forming regions in these galaxies, changing completely our picture of young stellar systems outside the Milky Way, and extending the extragalactic stellar IMF below the persisting threshold of a few solar masses. This review presents the recent developments in the investigation of the PMS stellar content of the Magellanic Clouds, with special focus on the limitations by single-epoch photometry that can only be circumvented by the detailed study of the observable behavior of these stars in the color–magnitude diagram. The achieved characterization of the low-mass PMS stars in the Magellanic Clouds allowed thus a more comprehensive understanding of the star formation process in our neighboring galaxies.

Published in Space Science Reviews Available from arXiv:1202.6534